# The Existence of MacWilliams-Type Identities for the Lee, Homogeneous and Subfield Metric

### v2

@@ 00 $1$ $1$ $1$ $1$
Given the finite field $\mathbb{F}_{27}=\mathbb{F}_{3}[\alpha]$ where $\alpha^{3}=\alpha+2$. Using the subfield-trace partition (as stated in (114)) we partition $\mathbb{F}_{27}$ into the following four blocks
(133) $I^{\lambda}_{0}$ =$=\left\{{0}\right\},\quad I^{\lambda}_{1}=\mathbb{F}_{27}^{\times}=\left\{{1,2}\right\},\quad I^{\lambda}_{2}=\left\{{\alpha,\alpha+1,\alpha+2,2\alpha,2\alpha+1,2\alpha+2}\right\},\quad\text{and}$
(134) $I^{\lambda}_{3}$ =$=\{\alpha^{2},\alpha^{2}+1,\alpha^{2}+2,\alpha^{2}+\alpha,\alpha^{2}+\alpha+1,\alpha^{2}+\alpha+2,\alpha^{2}+2\alpha,\alpha^{2}+2\alpha+1,\alpha^{2}+2\alpha+2,$
(135) $\qquad 2\alpha^{2},2\alpha^{2}+1,2\alpha^{2}+2,2\alpha^{2}+\alpha,2\alpha^{2}+\alpha+1,2\alpha^{2}+\alpha+2,2\alpha^{2}+2\alpha,2\alpha^{2}+2\alpha+1,2\alpha^{2}+2\alpha+2\}.$
@@
Consider the code $\mathcal{C}\subset\mathbb{F}_{27}^{3}$ generated by $G=\begin{pmatrix}1&\alpha&1\end{pmatrix}$. By computing the codewords of $\mathcal{C}$ we observe that they can be partitioned into six distinct subfield-trace decompositions, namely $(3,0,0,0),(0,2,1,0),(0,0,2,1),(0,0,1,2),(0,1,0,2)$ and $(0,0,0,3)$. Each of the six subfield-trace decompositions are represented by $1,2,6,4,2$ and $12$ codewords, respectively.
@@
The dual $\mathcal{C}^{\perp}$ of $\mathcal{C}$ has a generator matrix of the form
@@
We now verify that indeed $\frac{1}{|\mathcal{C}|}\sum_{\pi\in\mathbb{D}^{\lambda}_{p,r,1,n}(\rho)}K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=2$. We show detailed calculations of the Krawtchouk coefficient only for $\pi^{(0)}=(3,0,0,0)$ and $\pi^{(1)}=(0,2,1,0)$. Let us first assume, that $\pi=\pi^{(0)}$. Similar to the examples of the Lee and the homogeneous metric, we observe that
@@
For the first option $t\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)$ shown in Figure 2, we have $t_{01}=t_{11}=t_{21}=1$ and all other $t_{ij}$’s are equal to 00. The second option, i.e., $s\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)$, consists of $s_{11}=2$ and $s_{20}=1$ and $s_{ij}=0$ for all other indices. Hence,
(139) $K^{\lambda}_{\pi}(\rho)=\binom{2}{1,1}\binom{1}{1}(3-1)^{1+1}+\binom{2}{2}\binom{1}{1}(3-1)^{2}=12,\quad\text{and}\quad K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=24.$
@@
(140) $\frac{1}{|\mathcal{C}|}\sum_{\pi\in\mathbb{D}^{\lambda}_{p,r,1,n}(\rho)}K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=\frac{1}{27}\left(12+24+0\cdot 6+(-3)\cdot 4+(-3)\cdot 2+3\cdot 12\right)=\frac{1}{27}\cdot 54=2.$
Finally, we want to highlight at this point that the subfield-trace partition (see (117)) is indeed the coarsest partition that allows for the MacWilliams identities to exist and to recover the subfield weight enumerator.
4. Linear Programming Bound
Given a finite chain ring $\mathcal{R}$, an additive weight $\mathrm{wt}:\mathcal{R}\to\mathbb{Q}$ and an $\mathcal{R}$-linear code $\mathcal{C}$, recall the definition of the decomposition enumerator of $\mathcal{C}$
with respect to the weight $\operatorname{wt}$ (see (51), (81) and (118) for the Lee, homogeneous-unit and subfield-trace enumerator, respectively). Moreover, if $\left|\,{\mathbb{D}^{\operatorname{wt}}_{p,r,s,n}}\,\right|=D$, we denote by $\pi^{(0)}=(n,0,\dots,0)$ the decomposition of the all-zero vector and by $\pi^{(1)},\dots,\pi^{(D-1)}$ the other decompositions in $\mathbb{D}^{\operatorname{wt}}_{p,r,s,n}$.
The MacWilliams-type identities for the Lee decomposition enumerator, given in Corollary 26, allow us to straightforwardly derive an LP bound for any $\mathcal{R}$-linear code and any additive weight. Moreover, the MacWilliams-type identities given in Theorem 30 and Theorem 36 allow to derive an LP bound for the homogeneous and subfield metric respectively, which is less complex (i.e., that involves fewer variables).
From now on, $B_{p,r,s,n}(n,d_{\operatorname{wt}})$ denotes the maximum number of codewords of an $\mathcal{R}$-linear code of length $n$ and minimum $\mathrm{wt}$-distance $d_{\operatorname{wt}}$.
Theorem 38 (Linear Programming Bound).
For a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$ of minimum distance $d_{\operatorname{wt}}$ it holds that
(142) $B_{p,r,s,n}(n,d_{\operatorname{wt}})\leq\max\left(\sum_{i=0}^{D-1}\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})\right),$
  • $\mathcal{D}_{\pi^{(0)}}^{\operatorname{wt}}(\mathcal{C})=1$;
  • $\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})\geq 0$ for all $11\leq i\leq D-$ 1 and $\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})=0$ for all $i\in I$;
where $I=\{i\ \;|\;\ x\in\mathcal{R}^{n}\text{ such that }\pi(x)=\pi^{(i)}\text{ and }\mathrm{wt}(x)<d_{\operatorname{wt}}\}$.
If $\operatorname{wt}$ is an additive weight over $\mathcal{R}$ and $\mathcal{D}_{\pi}^{\operatorname{wt}}(\mathcal{C})$ is the Lee partition enumerator, the LP bound requires a large number of variables. Thus, the program cannot run efficiently and is not computationally feasible.
4.1. Lee metric
The Lee scheme, introduced in [6] and investigated in [4, 5, 41], provides a framework for studying the maximum size of a linear Lee-metric code. In the Lee-metric case the LP bound Theorem 38 can be refined as follows.
Through this section, let $\mathcal{R}=(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ and let $M$ be the maximum value of the Lee-weight in $\mathbb{Z}/{p^{s}}\mathbb{Z}$. Given a $\mathbb{Z}/{p^{s}}\mathbb{Z}$-linear code $\mathcal{C}$ of length $n$, let $c\in\mathcal{C}$ denote a codeword. In the Hamming metric, whenever we multiply each component of $c$ by a unit $r\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times}$, we obtain a codeword $r\cdot c$ of the same Hamming weight as $c$. This property does not hold in the Lee metric. However, the Lee decompositions of $c$ and $r\cdot c$ are related. Indeed, the Lee decomposition of $r\cdot c$ is a permutation of the Lee decomposition of $c$. That is, if $\pi^{\mathsf{L}}(c)=(\pi^{\mathsf{L}}_{0},\dots,\pi^{\mathsf{L}}_{M})$ is the Lee decomposition of $c$, then the Lee decomposition of $r\cdot c$ is given by
where $\sigma_{r}(i)=k$ such that $rk\equiv i\pmod{p^{s}}$.
Lemma 39 ([3, Lemma 1]).
Given a Lee decomposition $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$, a unit element $r\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times}$ and $\sigma_{r}$ a permutation on the set $\left\{{1,\ldots,M}\right\}$, we define $\sigma_{r}(\pi):=(\pi_{0},\pi_{{\sigma_{r}}(1)},\dots,\pi_{\sigma_{r}(M-1)})$. Then, for any linear code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$, it holds that $\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C})=\mathcal{D}_{\sigma(\pi)}^{\mathsf{L}}(\mathcal{C})$.
For any Lee decomposition $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$, we define the set
Since $\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C})=\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C})$ for every $\rho\in S(\pi)$, an additional constraint arises in the formulation of the LP problem for linear Lee-codes.
Let $B_{p,1,s,n}(n,d_{\mathsf{L}})$ denote the maximum number of codewords in a linear code of length $n$ and minimum Lee-distance $d_{\mathsf{L}}$ over $\mathbb{Z}/{p^{s}}\mathbb{Z}$.
Theorem 40 (Linear Programming Bound for Lee-metric Codes).
For a linear code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ of minimum Lee distance $d_{\mathsf{L}}$ it holds that
(145) $B_{p,1,s,n}(n,d_{\mathsf{L}})\leq\max\left(\sum_{i=0}^{D-1}\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})\right),$
  • $\mathcal{D}_{\pi^{(0)}}^{\mathsf{L}}(\mathcal{C})=1$;
  • $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})\geq 0$ for all $11\leq i\leq D-$ 1 and $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})=0$ for all $i\in I$;
  • and
where $I=\{i\ \;|\;\ x\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}\text{ such that }\pi(x)=\pi^{(i)}\text{ and }\operatorname{wt}_{\mathsf{L}}(x)<d_{\mathsf{L}}\}$.
The last constraint is derived from the fact that
5. Conclusion
In this paper we proved the existence of MacWilliams-type identities for any finite chain ring and any additive weight. In particular, this includes the homogeneous, the Lee and the subfield metric, where the classical MacWilliams identities fail. The main idea of the novel identities is to partition the ambient space into a finer partition, called Lee partition which is able to recover the weight enumerator of any additive weight. We then improved upon this result in the case of the homogeneous and the subfield metric, by presenting a coarser partition than the Lee partition, but finer than the weight partition, for which MacWilliams-type identities hold. The presented identities can be used to formulate LP bounds in each of the metrics.

### v3

@@ A.1. Lee Weight Partition
Consider the code $\mathcal{C}\subseteq(\mathbb{Z}/{9}\mathbb{Z})^{3}$ generated by $G=\begin{pmatrix}3&2&8\end{pmatrix}$. That is,
(85) $\mathcal{C}=\big\{(0,0,0),$ $(0,3,3),(0,6,6),(3,2,8),(6,7,1),(6,4,7),(3,5,2),(3,8,5),(6,1,4)\big\}.$
$\pi^{(0)}=(3,0,0,0,0),\ \pi^{(1)}=(1,0,0,2,0),\ \pi^{(2)}=(0,1,1,1,0),$
$\pi^{(3)}=(0,0,1,1,1)\ \text{ and }\ \pi^{(4)}=(0,1,0,1,1).$
(86) $G^{\perp}=\begin{pmatrix}1&0&3\\ 0&1&2\end{pmatrix}$.
(87) $\prod_{i=0}^{M}\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}\prod_{j=1}^{M}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}$ $=\binom{\pi_{0}}{t_{00},\ldots,t_{0M}}\left(1+1\right)^{t_{03}}$
(88) $=\binom{3}{2,1}(1+1)$
(89) $=6.$
Consider now the partition $\pi=\pi^{(1)}=(1,0,0,2,0)$. There are two compositions $t=(t_{0},\ldots,t_{M}),s=(s_{0},\ldots,s_{M})\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)$, given by
(90) $t_{0}=(1,0,0,0,0),\;t_{3}=(1,0,0,1,0)\text{ and }t_{1}=t_{2}=t_{4}=(0,\ldots,0),$
(91) $s_{0}=(0,0,0,1,0),\;s_{3}=(2,0,0,0,0)\text{ and }s_{1}=s_{2}=s_{4}=(0,\ldots,0).$
Hence, we get
(92) $\sum_{t\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)}\left(\prod_{i=0}^{4}\binom{\pi_{i}}{t_{i0},\ldots,t_{i4}}\prod_{j=1}^{4}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}\right)$ $=\left(\binom{\pi_{0}}{t_{00},\ldots,t_{04}}\prod_{j=1}^{4}\left(\xi^{0}+\xi^{0}\right)^{t_{0j}}\binom{\pi_{3}}{t_{30},\ldots,t_{34}}\prod_{j=1}^{M}\left(\xi^{-3j}+\xi^{3j}\right)^{t_{3j}}\right$.
(93) $\quad\qquad+\left.\binom{\pi_{0}}{s_{00},\ldots,s_{04}}\prod_{j=1}^{4}\left(\xi^{0}+\xi^{0}\right)^{s_{0j}}\binom{\pi_{3}}{s_{30},\ldots,s_{34}}\prod_{j=1}^{M}\left(\xi^{-3j}+\xi^{3j}\right)^{s_{3j}}\right)$
(94) $=\left(\binom{1}{1}\binom{2}{1,1}(\xi^{0}+\xi^{0})^{1}\binom{1}{1}\binom{2}{2}(\xi^{0}+\xi^{0})^{1}\right)$
(95) $=6.$
For $\pi\in\left\{{\pi^{(2)},\pi^{(3)},\pi^{(4)}}\right\}$ the computations are similar and yield in each case
since $\sum_{i\in\left\langle{3}\right\rangle}\xi^{i}=0$. Putting the expressions given in (87), (92) and (96) in the right-hand side of (41) yields
(97) $\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{L}}_{n}}K^{\mathsf{L}}_{\pi}(\rho)\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C})=\frac{1}{9}\left(6\cdot 1+6\cdot 2+0\cdot 2+0\cdot 2+0\cdot 2\right)=2,$
A.2. Homogeneous Weight-Unit Partition
Let us now consider an example for the MacWilliams-type identities for the homogeneous weight-unit partition stated in Theorem 21.
Example 39.
Given $q=2$ and $s=4$, consider the code $\mathcal{C}\subset(\mathbb{Z}/{16}\mathbb{Z})^{3}$ generated by $G=\begin{pmatrix}8&1&8\end{pmatrix}$. Notice, that there are four distinct homogeneous weight-unit decompositions occurring in $\mathcal{C}$, namely,
$\pi^{(0)}=(3,0,0,0),\quad\pi^{(1)}=(2,0,0,1),\quad\pi^{(2)}=(2,0,1,0),\quad\pi^{(3)}=(0,1,2,0),$
(98) $\mathcal{D}_{\pi^{(0)}}^{\mathsf{hom}}(\mathcal{C})=1,\quad\mathcal{D}_{\pi^{(1)}}^{\mathsf{hom}}(\mathcal{C})=6,\quad\mathcal{D}_{\pi^{(2)}}^{\mathsf{hom}}(\mathcal{C})=1,\quad\mathcal{D}_{\pi^{(3)}}^{\mathsf{hom}}(\mathcal{C})=8.$
Note that the dual code of $\mathcal{C}$, $\mathcal{C}^{\perp}$, is generated by
(99) $G^{\perp}=\begin{pmatrix}1&8&0\\ 0&8&1\end{pmatrix}$.
Let $\rho=(1,1,1,0)\in\mathbb{D}_{3}^{\mathsf{hom}}$ be a homogeneous weight-unit decomposition in $(\mathbb{Z}/{16}\mathbb{Z})^{3}$. By calculating the codewords of $\mathcal{C}^{\perp}$, we observe that $\mathcal{D}_{\rho}^{\mathsf{hom}}(\mathcal{C}^{\perp})=16$. For every $\pi\in\left\{{\pi^{(0)},\pi^{(1)},\pi^{(2)},\pi^{(3)}}\right\}$ let us now compute the Krawtchouk coefficient $K^{\mathsf{hom}}_{\pi}(\rho)$ as stated in Theorem 28.
(100) $\mathrm{Comp}^{\mathsf{hom}}_{\pi^{(0)}}(\rho)=\Big\{(\underbrace{1,1,1,0}_{t_{Z}},\underbrace{0,0,0,0}_{t_{U}},\underbrace{0,0,0,0}_{t_{S}},\underbrace{0,0,0,0}_{t_{R}})\Big\}.$
Hence, we obtain
(101) $K^{\mathsf{hom}}_{\pi^{(0)}}(\rho)$ $=\prod_{i\in\left\{{Z,U,S,R}\right\}}\binom{\pi_{i}}{t_{iZ},\ldots,t_{iR}}=\binom{\pi_{Z}}{t_{Z}}(q^{s-1}(q-1))^{t_{ZU}}$
(102) $=\binom{3}{1,1,1,0}8=48.$
Now, let $\pi=\pi^{(1)}$. Note that the set of compositions $\mathrm{Comp}^{\mathsf{hom}}_{\pi}(\rho)$ of $\rho$ with respect to $\pi$ consists of only three elements, i.e.,
(103) $\mathrm{Comp}^{\mathsf{hom}}_{\pi^{(1)}}(\rho)=\Big\{$ $t=(\underbrace{1,1,0,0}_{t_{Z}},\underbrace{0,0,0,0}_{t_{U}},\underbrace{0,0,0,0}_{t_{S}},\underbrace{0,0,1,0}_{t_{R}}),$
(104) $s=(\underbrace{1,0,1,0}_{s_{Z}},\underbrace{0,0,0,0}_{s_{U}},\underbrace{0,0,0,0}_{s_{S}},\underbrace{0,1,0,0}_{s_{R}})$
(105) $v=(\underbrace{0,1,1,0}_{v_{Z}},\underbrace{0,0,0,0}_{v_{U}},\underbrace{0,0,0,0}_{v_{S}},\underbrace{1,0,0,0}_{s_{R}})\Big\}.$
Recalling that q1=1q-1=1 and omitting all the powers of $1$, we obtain
(106) $K^{\mathsf{hom}}_{\pi^{(1)}}(\rho)$ $=\binom{\pi_{Z}}{t_{Z}}\binom{\pi_{R}}{t_{R}}\left(q^{s-1}\right)^{t_{ZU}}+\binom{\pi_{Z}}{v_{Z}}\binom{\pi_{R}}{v_{R}}(q^{s-1})^{v_{ZU}}$
(107) $=\binom{2}{1,1,0,0}\binom{1}{0,0,1,0}8+\binom{2}{0,1,1,0}\binom{1}{1,0,0,0}8$
(108) $=32.$
Moreover, if $\pi=\pi^{(2)}$, we have
(109) $\mathrm{Comp}^{\mathsf{hom}}_{\pi^{(2)}}(\rho)=\Big\{$ $t=(\underbrace{1,1,0,0}_{t_{Z}},\underbrace{0,0,0,0}_{t_{U}},\underbrace{0,0,1,0}_{t_{S}},\underbrace{0,0,0,0}_{t_{R}}),$
(110) $s=(\underbrace{1,0,1,0}_{s_{Z}},\underbrace{0,0,0,0}_{s_{U}},\underbrace{0,1,0,0}_{s_{S}},\underbrace{0,0,0,0}_{s_{R}})$
(111) $v=(\underbrace{0,1,1,0}_{v_{Z}},\underbrace{0,0,0,0}_{v_{U}},\underbrace{1,0,0,0}_{v_{S}},\underbrace{0,0,0,0}_{s_{R}})\Big\}.$
It follows that
(112) $K^{\mathsf{hom}}_{\pi^{(2)}}(\rho)$ $=\binom{\pi_{Z}}{t_{Z}}\binom{\pi_{S}}{t_{S}}\left(q^{s-1}\right)^{t_{ZU}}+\binom{\pi_{Z}}{s_{Z}}\binom{\pi_{S}}{s_{S}}(-1)^{s_{SU}}+\binom{\pi_{Z}}{v_{Z}}\binom{\pi_{S}}{v_{S}}(q^{s-1})^{v_{ZU}}$
(113) $=\binom{2}{1,1,0,0}\binom{1}{0,0,1,0}8+\binom{2}{1,0,1,0}\binom{1}{0,1,0,0}(-8)+\binom{2}{0,1,1,0}\binom{1}{1,0,0,0}8$
(114) $=16.$
Lastly, for $\pi^{(3)}$, we have
(115) $\mathrm{Comp}^{\mathsf{hom}}_{\pi^{(2)}}(\rho)=\Big\{$ $t=(\underbrace{0,0,0,0}_{t_{Z}},\underbrace{1,0,0,0}_{t_{U}},\underbrace{0,1,1,0}_{t_{S}},\underbrace{0,0,0,0}_{t_{R}}),$
(116) $s=(\underbrace{0,0,0,0}_{s_{Z}},\underbrace{0,1,0,0}_{s_{U}},\underbrace{1,0,1,0}_{s_{S}},\underbrace{0,0,0,0}_{s_{R}})$
(117) $v=(\underbrace{0,0,0,0}_{v_{Z}},\underbrace{0,0,1,0}_{v_{U}},\underbrace{1,1,0,0}_{v_{S}},\underbrace{0,0,0,0}_{s_{R}})\Big\},$
which yields to
(118) $K^{\mathsf{hom}}_{\pi^{(2)}}(\rho)$ $=\binom{\pi_{U}}{t_{U}}\binom{\pi_{S}}{t_{S}}\left(-q^{s-1}\right)^{t_{SU}}+\binom{\pi_{U}}{s_{U}}\binom{\pi_{S}}{s_{S}}0^{s_{UU}}+\binom{\pi_{U}}{v_{U}}\binom{\pi_{S}}{v_{S}}(-1)^{v_{US}}(-q^{s-1})^{v_{ZU}}$
(119) $=\binom{1}{1,0,0,0}\binom{2}{0,1,1,0}(-8)+\binom{1}{0,0,1,0}\binom{2}{1,1,0,0}8$
(120) $=0.$
Finally, inserting these values in Equation (62), yields
(121) $\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{L}}_{n}}K^{\mathsf{L}}_{\pi}(\rho)\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C})=\frac{1}{16}(48\cdot 1+6\cdot 32+1\cdot 16+0\cdot 8)=16,$
which corresponds with $\mathcal{D}_{\rho}^{\mathsf{hom}}(\mathcal{C}^{\perp})$.
A.3. Subfield-Trace Partition
Lastly, we give an example to show the calculations of the MacWilliams-type identities of Corollary 33.
Example 40.
@@ A.3. Subfield-Trace Partition
Given the finite field $\mathbb{F}_{27}=\mathbb{F}_{3}[\alpha]$ where $\alpha^{3}=\alpha+2$. Using the subfield-trace partition (as stated in (LABEL:eq:subfield-trace_Fpr)) we partition $\mathbb{F}_{27}$ into the following four blocks
$I^{\lambda}_{0}=\left\{{0}\right\},\quad I^{\lambda}_{1}=\mathbb{F}_{27}^{\times}=\left\{{1,2}\right\},\quad I^{\lambda}_{2}=\left\{{\alpha,\alpha+1,\alpha+2,2\alpha,2\alpha+1,2\alpha+2}\right\},\quad$
$\text{and }\quad I^{\lambda}_{3}=\mathbb{F}_{27}\setminus\left(I^{\lambda}_{0}\cup I^{\lambda}_{1}\cup I^{\lambda}_{2}\right)$
@@
Consider the code $\mathcal{C}\subset\mathbb{F}_{27}^{3}$ generated by $G=\begin{pmatrix}1&\alpha&1\end{pmatrix}$. By computing the codewords of $\mathcal{C}$ we observe that they can be partitioned into six distinct subfield-trace decompositions, namely
$(3,0,0,0),(0,2,1,0),(0,0,2,1),(0,0,1,2),(0,1,0,2)\text{ and }(0,0,0,3).$
@@
Each of the six subfield-trace decompositions are represented by $1,2,6,4,2$ and $12$ codewords, respectively. The dual $\mathcal{C}^{\perp}$ of $\mathcal{C}$ has a generator matrix of the form
@@
We now verify that indeed $\frac{1}{|\mathcal{C}|}\sum_{\pi\in\mathbb{D}^{\lambda}_{n}(\rho)}K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=2$. We show detailed calculations of the Krawtchouk coefficient only for $\pi^{(0)}=(3,0,0,0)$ and $\pi^{(1)}=(0,2,1,0)$. Let us first assume, that $\pi=\pi^{(0)}$. Similar to the examples of the Lee and the homogeneous metric, we observe that
@@
For the first option $t\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)$ shown in Figure 2, we have $t_{01}=t_{11}=t_{21}=1$ and all other $t_{ij}$’s are equal to 0. The second option, i.e., $s\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)$, consists of $s_{11}=2$ and $s_{20}=1$ and $s_{ij}=0$ for all other indices. Hence,
$K^{\lambda}_{\pi}(\rho)=\binom{2}{1,1}\binom{1}{1}(3-1)^{1+1}+\binom{2}{2}\binom{1}{1}(3-1)^{2}=12\quad\text{and}\quad K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=24.$
@@
(125) $\frac{1}{|\mathcal{C}|}\sum_{\pi\in\mathbb{D}^{\lambda}_{n}(\rho)}K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C})=\frac{1}{27}\left(12+24+0-12-6+36\right)=\frac{1}{27}\cdot 54=2.$